OpenStudy (deadshot):

What are the zeros of f(x) = x^2 - 5x -24?

OpenStudy (akashdeepdeb):

Split the middle term! :)

OpenStudy (akashdeepdeb):

Getting me?

OpenStudy (deadshot):

not quite, do you mean split like x^2 + 1x +1 = (x+1)^2

OpenStudy (deadshot):

\[x^2-5x-24=(x+3)(x-8)\] right?

OpenStudy (akashdeepdeb):

EXACTLY! :D

OpenStudy (deadshot):

so then the zeroes are 3 and -8, right?

OpenStudy (akashdeepdeb):

Not quite :D If (x+3)(x-8) = 0 then (x+3) = 0/(x-8) = 0 So x = -3 And same for (x_8)!! :D Understood? :)

OpenStudy (deadshot):

yes, thanks!

OpenStudy (akashdeepdeb):

:)